What is the value of Q when the SrCrO4 solution contains 2.00×10−2 M Sr2+ and 1.50×10−3M CrO42−? Calculate the value of Q.
Answer #1

SrCrO4 <--> Sr+2 and CrO4-2

Q = [Sr+2][CrO4-2]

Q = (2*10^-2)(1.5*10^-3) = 3*10^-5

Q = 3*10^-5